OpenStudy (anonymous):

Solve. (x + 2)^2 = 1

OpenStudy (anonymous):

completing the square

OpenStudy (ajprincess):

take square root both sides \(\sqrt{(x+2)^2}=\pm\sqrt{1^2}\) \(x+2=~or~x+2=-1\)

OpenStudy (anonymous):

(x+2)^2=1 x^2+4x+4=1 x^2+4x+3=0 x^2+x+3x+3=0 x(x+1)+3(x+1)=0 (x+1)(x+3)=0 So, x=-1 or x= -3

OpenStudy (ajprincess):

* x+2=1 or x+2=-1

OpenStudy (anonymous):

x=1 or x=-1 ?

OpenStudy (ajprincess):

x+2=1 x=1-2 =-1 x+2=-1 x=-2-1=?